Solve equation.
5/6x -1/5x + 4 = 7/10x
Having difficultly with fractions

Answer by checkley79(3341)

5/6x -1/5x + 4 = 7/10x
5X/6-X/5+4=7X/10
5X/6-X/5-7X/10=-4
(5X*5-6X-7X*3)/30=-4
(25X-6X-21X)/30=-4
-2X/30=-4
-2X=-4*30
-2X=-120
X=-120/-2
X=60 ANS.
PROOF:
(5/6)*60-(1/5)*60+4=(7/10)60
300/6-60/5+4=420/10
50-12+4=42
54-12=42
42=42